The question states:
Each set of ordered pairs is for a direct variation. Find each missing value.
I know the answer which is 6 because it says so in the back of the book.
(4,8) and (3,y)
Can anyone explain the steps in the process to me. I have a huge test tomorrow.

Answer by stanbon(75887)

Each set of ordered pairs is for a direct variation. Find each missing value.
I know the answer which is 6 because it says so in the back of the book.
(4,8) and (3,y)
----------------
Direct variation: y=kx
Using (4,8), solve for "k": 8=k*4; k=2
So, y=2x
Using (3,y), solve for "y": y=2*3; y=6
Cheers,
Stan H.